Find the range space and a basis for the
kernel of the linear transformation
T : R4 ->R4 defined by
T(x1, x2, x3, x4) = (x1 - x2, x2 - x3, x3 - x4, x4 - x1).
1
Expert's answer
2020-12-02T11:06:15-0500

"1.\text{ Kernel of a transformation}.\\
T: \mathbb{R}^4\longrightarrow \mathbb{R}^4\\
T(x_1,x_2,x_3,x_4)=(x_1-x_2,x_2-x_3,x_3-x_4,x_4-x_1)\\
\begin{cases}
x_1-x_2=0\\
x_2-x_3=0\\
x_3-x_4=0\\
x_4-x_1=0
\end{cases}\\
\begin{pmatrix}
1&-1&0&0\\
0&1&-1&0\\
0&0&1&-1\\
-1&0&0&1
\end{pmatrix}\\
\text{We add the first row to the fourth row:}\\
\begin{pmatrix}
1&-1&0&0\\
0&1&-1&0\\
0&0&1&-1\\
0&-1&0&1
\end{pmatrix}\\
\text{We add the second row to the fourth row:}\\
\begin{pmatrix}
1&-1&0&0\\
0&1&-1&0\\
0&0&1&-1\\
0&0&-1&1
\end{pmatrix}\\
\text{We add the third row to the fourth row:}\\
\begin{pmatrix}
1&-1&0&0\\
0&1&-1&0\\
0&0&1&-1\\
0&0&0&0
\end{pmatrix}\\
\begin{cases}
x_1-x_2=0\\
x_2-x_3=0\\
x_3-x_4=0
\end{cases}\\
\begin{cases}
x_1=x_2\\
x_2=x_3\\
x_3=x_4
\end{cases}\\
x_1=x_2=x_3=x_4\\
\text{ker}(T)=\{(x_1,x_2,x_3,x_4)\in\mathbb{R}^4|x_1=x_2=x_3=x_4\}.\\
\text{Hence } \{(1,1,1,1)\}\text{ is a basis of ker(T)}.\\
2.\text{ Range space}.\\
\text{Let } (a,b,c,d)\in \mathbb{R}^4.\\
\mathbb{R}^4\text{ is a codomain of T}.\\
\text{Then:}\\
\begin{cases}
x_1-x_2=a\\
x_2-x_3=b\\
x_3-x_4=c\\
x_4-x_1=d
\end{cases}\\
\text{We have shown above that the rank of the coefficient matrix equals 3}.\\
\text{Using Rouché–Capelli theorem:}\\
\text{rank}
\begin{pmatrix}
1&-1&0&0&a\\
0&1&-1&0&b\\
0&0&1&-1&c\\
-1&0&0&1&d
\end{pmatrix}=3\\
\text{So the range space of T consists of all the vectors }(a,b,c,d)\in\mathbb{R}^4\\
\text{such that the last equality is true.} \\
\begin{pmatrix}
1&-1&0&0&a\\
0&1&-1&0&b\\
0&0&1&-1&c\\
-1&0&0&1&d
\end{pmatrix}\sim
\begin{pmatrix}
1&-1&0&0&a\\
0&1&-1&0&b\\
0&0&1&-1&c\\
0&-1&0&1&a+d
\end{pmatrix}\sim
\begin{pmatrix}
1&-1&0&0&a\\
0&1&-1&0&b\\
0&0&1&-1&c\\
0&0&-1&1&a+b+d
\end{pmatrix}\sim
\begin{pmatrix}
1&-1&0&0&a\\
0&1&-1&0&b\\
0&0&1&-1&c\\
0&0&0&0&a+b+c+d
\end{pmatrix}\\
\text{Then } a+b+c+d=0.\\
\text{The range space of T consists of all the vectors }(a,b,c,d)\in\mathbb{R}^4\\
\text{such that }a+b+c+d=0."